# Conditional Brownian Bridge Diffusion Model for VHR SAR to Optical Image Translation

### v3

@@ Abstract
Synthetic Aperture Radar (SAR) imaging technology provides the unique advantage of being able to collect data regardless of weather conditions and time. However, SAR images exhibit complex backscatter patterns and speckle noise, which necessitate expertise for interpretation. Research on translating SAR images into optical-like representations has been conducted to aid the interpretation of SAR data. Nevertheless, existing studies have predominantly utilized low-resolution satellite imagery datasets and have largely been based on Generative Adversarial Network (GAN) which are known for their training instability and low fidelity. To overcome these limitations of low-resolution data usage and GAN-based approaches, this paper introduces a conditional image-to-image translation approach based on Brownian Bridge Diffusion Model (BBDM). We conducted comprehensive experiments on the MSAW dataset, a paired SAR and optical images collection of 0.5m Very-High-Resolution (VHR). The experimental results indicate that our method surpasses both the Conditional Diffusion Models (CDMs) and the GAN-based models in diverse perceptual quality metrics.
@@ I Introduction
Yet, even with this inherent advantage of utilizing these low-resolution images, existing GAN-based approaches for translating SAR-to-optical have struggled to achieve practical performance, facing issues like training instability, mode collapse, and geometric loss with complex scenes [1, 2, 3], [9]. Only a few recent studies [4, 5] have explored Conditional Diffusion Models (CDMs) to overcome these limitations of GAN-based models. CDMs currently dominate the image synthesis field [10] including image-to-image translation. Still, CDMs suffer from difficulties with limited model generalization despite their potential. They lack robust theoretical foundations to ensure that the outcome accurately represents the intended conditional distribution [11]. These models often experience performance degradation when translating between significantly disparate domains.
@@
Basically, the original BBDM performs the diffusion process in a compressed latent space to operate it more efficiently by leveraging compact yet informative latent representations. Inspired by CDMs [10], we adjusted the BBDM framework for VHR SAR to optical image translation. Our approach incorporates bilinearly interpolated information from the pixel space as a condition. This ensures that the model receives spatial information to guide the translation process. This combination of latent representations and conditioned information strikes a balance between computational efficiency and the preservation of essential spatial details from the SAR image.
@@
The main contributions in this paper include the following:
@@
2) By incorporating spatially interpolated information from the pixel space as a condition into BBDM, we achieved substantial improvements in both structural fidelity and visual quality for VHR SAR to optical imagery translation.
@@ II-C Conditional BBDM for SAR to optical
TABLE I: PERFORMANCE COMPARISON of DIFFERENT MODELS for VHR SAR to OPTICAL IMAGE TRANSLATION
Methods LPIPS-ALEX↓ LPIPS-VGG↓ LPIPS-SQUEEZE↓ FID↓
Pix2pix 0.5298 0.6195 0.3690 241.58
CycleGAN 0.5395 0.6429 0.4254 222.95
Conditional LDM 0.6268 0.6333 0.4477 256.17
BBDM 0.4946 0.5952 0.3599 197.54
Conditional BBDM (Ours) 0.4579 0.5744 0.3286 177.79
@@ III-B Evaluation Metrics
We used several metrics to assess image translation performance of our proposed method and to compare it with conditional LDM and GAN-based models. Three Learned Perceptual Image Patch Similarity (LPIPS) [15] scores are selected to evaluate how effectively the translated SAR imagery improves SAR image interpretability. These scores, based on three different neural networks (AlexNet, VGG-16, and SqueezeNet), are widely used to measure perceptual quality. We also used Fréchet Inception Distance (FID) [16] to evaluate distribution similarity between generated and authentic optical images. Lower scores indicate better performance for both metrics.
@@ III-C Implementation Details
We adopted a pre-trained VQ-VAE that reduces the pixel space dimensions by a factor of 4 in both the height and width directions. We trained the proposed model for 48 epochs using an NVIDIA A6000 GPU, with a minibatch size of 1 and Adam optimizer of 1e-4 learning rate. Data augmentation was implemented using horizontal flips during training. To ensure fair comparison, we applied same optimizer and augmentation techniques when training the conditional LDM and GAN-based models. These comparative models were trained using their default minibatch size and learning rates, which are widely adopted in the existing literature. The conditional LDM was fine-tuned from the pre-trained model by [13]. For evaluation, we selected the checkpoint of each model that recorded the minimal loss on the validation dataset.
@@ III-D1 Quantitative Results
Table I presents the quantitative comparison between ground truth optical images and translated optical-like images. The top score is marked in red and the second-highest is marked in green. Our proposed model significantly outperforms other models in three different LPIPS scores and FID. The consistently superior LPIPS scores across all three measures demonstrate the robustness of our model’s perceptual quality. This consistent performance highlights the reliability of our proposed method as an effective SAR-to-optical translation framework. The best FID score suggests our model most accurately reconstructs the target distribution. These strong results on deep learning-based metrics highlight the potential of the proposed model’s generated optical-like images for deep learning-based downstream applications, such as synthetic pretraining.
@@ IV Conclusion
In this paper, we have proposed a new conditional BBDM framework for VHR SAR to optical image translation. Our proposed method achieved high quality and cost-effective translation for large VHR images through its integration of conditional information with Brownian bridge mapping in latent space. Quantitative and qualitative assessments indicate that proposed method generates impressive optical-like images and reveals its practical potential for SAR imagery interpretation. As an alternative to GAN-based approaches, our work contributes to improving SAR image interpretation and narrowing the gap between SAR and optical remote sensing modalities. This development potentially leads to improved applications.

### v4

@@ Abstract
Synthetic Aperture Radar (SAR) imaging technology provides the unique advantage of being able to collect data regardless of weather conditions and time. However, SAR images exhibit complex backscatter patterns and speckle noise, which necessitate expertise for interpretation. Research on translating SAR images into optical-like representations has been conducted to aid the interpretation of SAR data. Nevertheless, existing studies have predominantly utilized low-resolution satellite imagery datasets and have largely been based on Generative Adversarial Network (GAN) which are known for their training instability and low fidelity. To overcome these limitations of low-resolution data usage and GAN-based approaches, this letter introduces a conditional image-to-image translation approach based on Brownian Bridge Diffusion Model (BBDM). We conducted comprehensive experiments on the MSAW dataset, a paired SAR and optical images collection of 0.5m Very-High-Resolution (VHR). The experimental results indicate that our method surpasses both the Conditional Diffusion Models (CDMs) and the GAN-based models in diverse perceptual quality metrics.
@@ I Introduction
Yet, even with this inherent advantage of utilizing these low-resolution images, existing GAN-based approaches for translating SAR-to-optical have struggled to achieve practical performance, facing issues like training instability, mode collapse, and geometric loss with complex scenes [1, 2, 3], [9]. Only a few recent studies [4, 5] have explored Conditional Diffusion Models (CDMs) to overcome these limitations of GAN-based models. More recently, Shi et al. [10] proposed a brain-inspired approach that enhances CDMs by incorporating self-attention mechanisms and long-skip connections to improve feature extraction capabilities. CDMs currently dominate the image synthesis field [11] including image-to-image translation. Still, CDMs suffer from difficulties with limited model generalization despite their potential. They lack robust theoretical foundations to ensure that the outcome accurately represents the intended conditional distribution [12]. These models often experience performance degradation when translating between significantly disparate domains.
@@
Basically, the original BBDM performs the diffusion process in a compressed latent space to operate it more efficiently by leveraging compact yet informative latent representations. While this approach offers advantages over GANs and conventional diffusion models through stable training with its Brownian Bridge mapping incorporated gradual denoising process, it lacks mechanisms for domain-specific guidance. Inspired by CDMs [11], we adjusted the BBDM framework for VHR SAR to optical image translation. Our approach incorporates bilinearly interpolated information from the pixel space as a condition, which fundamentally differs from the original BBDM’s unconditioned approach. This ensures that the model receives spatial information to guide the translation process. This combination of latent representations and conditioned information not only maintains the stable training advantages of the original BBDM over GANs, but also overcomes the original BBDM’s limitations by striking a balance between computational efficiency and the preservation of essential spatial details from the SAR image.
@@
The main contributions in this letter include the following:
@@
2) In contrast to the original BBDM’s unconditioned approach, by incorporating spatially interpolated information from the pixel space as a condition into BBDM, we achieved substantial improvements in both structural fidelity and visual quality for VHR SAR to optical imagery translation.
@@ II-C Conditional BBDM for SAR to optical
TABLE I: Performance Comparison of Different Models for VHR SAR to Optical Image Translation. The Directional Arrows adjacent to each Evaluation Metric Indicate the Metric’s Correlation with Better Quality. The Highest score is Marked in Red and the Second-Highest score is Marked in Green
Methods LPIPS-ALEX↓ LPIPS-VGG↓ LPIPS-SQUEEZE↓ FID↓ SAM↓ CHD↓ CW-SSIM↑ FSIMc↑
Pix2pix 0.5298 0.6195 0.3690 241.58 0.0734 0.2865 0.4100 0.6532
CycleGAN 0.5395 0.6429 0.4254 222.95 0.0903 0.2961 0.4058 0.6396
Conditional LDM 0.6268 0.6333 0.4477 256.17 0.0882 0.5317 0.3568 0.6250
BBDM 0.4946 0.5952 0.3599 197.54 0.0728 0.2785 0.3905 0.6462
Conditional BBDM (Ours) 0.4579 0.5744 0.3286 177.79 0.0709 0.2411 0.4138 0.6595
@@ III-B Evaluation Metrics
We used several metrics to assess image translation performance of our proposed method and to compare it with conditional LDM and GAN-based models. We utilized three different Learned Perceptual Image Patch Similarity (LPIPS) scores and Fréchet Inception Distance (FID), to evaluate perceptual similarity and distribution similarity between generated and authentic optical images, respectively. We further included Spectral Angle Mapper (SAM) and Color Histogram Distance (CHD) for spectral consistency evaluation. SAM measures the angular difference of spectral patterns and CHD measures the global difference of color histograms between two images. Additionally, to evaluate spatial quality, we utilized Complex Wavelet Structural Similarity (CW-SSIM) index, which measures structural coherence in the complex wavelet domain, and Chrominance Feature Similarity (FSIMc) index, which effectively captures key structural features such as edges and textures in the image[16].
@@ III-C Implementation Details
We adopted a pre-trained Vector Quantized-Variational Auto Encoder(VQ-VAE) [17], a generative model that learns discrete latent representations by using a codebook to effectively reduce the pixel space dimensions by a factor of 4 in both the height and width directions. We trained the proposed model for 48 epochs using an NVIDIA A6000 GPU, with a minibatch size of 1 and Adam optimizer of 1e-4 learning rate. Data augmentation was implemented using horizontal flips during training. To ensure fair comparison, we applied same optimizer and augmentation techniques when training the conditional LDM and GAN-based models. These comparative models were trained using their default minibatch size and learning rates, which are widely adopted in the existing literature. The conditional LDM was fine-tuned from the pre-trained model by [14]. For evaluation, we selected the checkpoint of each model that recorded the minimal loss on the validation dataset.
@@ III-D1 Quantitative Results
Table I presents the quantitative comparison between ground truth optical images and translated optical-like images. Our proposed model significantly outperforms other models in all metrics. The consistently superior LPIPS scores demonstrate our model’s robustly high perceptual quality. The best FID score suggests our model most accurately reconstructs the target distribution. Beyond perceptual metrics, our method also shows the best quality in spectral and structural consistency. The lowest SAM and CHD values indicate our model’s capability to maintain spectral characteristics and color distributions of authentic optical images. The highest CW-SSIM value confirms our approach’s ability to maintain structural coherence. The best FSIMc score demonstrates that our method effectively maintains texture details and local structural features. Our model’s superior performance in deep learning-based metrics like LPIPS and FID demonstrates its potential for deep learning-based downstream applications such as synthetic pretraining. Simultaneously, the best structural and spectral similarity metrics suggest that our approach can effectively assist SAR image interpretation by translating complicated SAR data into visually intuitive optical images.
@@ IV Conclusion
In this letter, we have proposed a new conditional BBDM framework for VHR SAR to optical image translation. Unlike the original BBDM which operates solely in latent space, our approach combines the benefits of Brownian Bridge mapping in latent space with explicit spatial conditioning from the pixel domain. With this integration, our proposed method achieved high quality and cost-effective translation for large VHR images. Quantitative and qualitative assessments indicate that proposed method generates impressive optical-like images and reveals its practical potential for SAR imagery interpretation. As an alternative to GAN-based approaches, our work contributes to improving SAR image interpretation and narrowing the gap between SAR and optical remote sensing modalities. This development potentially leads to improved applications. However, our study has limitations that warrant further investigation. The current research lacks comprehensive ablation studies exploring different spatial feature extraction methods and conditioning mechanisms. Therefore, in future work, we aim to supplement these studies and conduct architectural refinements for further improvement.